# Elliptic Quantum Toroidal Algebra $U_{t_{1},t_{2},p}({\mathfrak{gl}}_{1,tor}),$ Vertex Operators and $L$-operators Hitoshi Konno and Andrey Smirnov

### v1

@@ Acknowledgements
$\Phi^{D}(u):{\cal F}^{(1,N)}_{v}\to{\cal F}^{(0,1)}_{u}\widetilde{\otimes}{\cal F}^{(1,N-1)}_{-v/u},$
$\Psi^{D*}(u)\ :\ {\cal F}^{(1,N)}_{v}\widetilde{\otimes}{\cal F}^{(0,1)}_{u}\to{\cal F}^{(1,N+1)}_{-vu}$
@@
$x\Psi^{D*}(v)=\Psi^{D*}(v)\Delta^{D}(x)\qquad\qquad\forall x\in{\cal U}.$ (A.2)
@@
$\Phi^{D}(u)\xi=\sum_{\lambda\in{\cal P}^{+}}{|\lambda\rangle}^{\prime}_{u}\widetilde{\otimes}\Phi^{D}_{\lambda}(u)\xi,$ (A.3)
$\Psi^{D*}_{\lambda}(u)\xi=\Psi^{D*}(u)\left(\xi\widetilde{\otimes}{|\lambda\rangle}_{u}\right)\quad\forall\xi\in{\cal F}^{(1,N)}_{v},$ (A.4)
@@
${\Phi}^{D}_{\lambda}(u)=\frac{t_{1}^{-n(\lambda^{\prime})-|\lambda|}N_{\lambda}(p)}{c_{\lambda}}\ :{\widetilde{\Phi}}_{\emptyset}(u)\prod_{\square\in\lambda}{x}^{-}(\varphi^{\lambda}_{\square}u):(-u)^{-\Lambda_{0}}e^{-\Lambda_{0}^{\perp}},$ (A.8)
${\Psi}^{D*}_{\lambda}(u)=\frac{t_{1}^{-n(\lambda^{\prime})}}{c_{\lambda}N_{\lambda}^{\prime}(p^{*})}\ :\widetilde{\Psi}^{*}_{\emptyset}(u)\prod_{\square\in\lambda}{x}^{+}(\varphi^{\lambda}_{\square}u):(-u)^{\Lambda_{0}}e^{\Lambda_{0}^{\perp}},$ (A.9)
@@
$N_{\lambda}(p)=\prod_{\Box\in\lambda}\frac{(pt_{1}^{-a(\Box)-1}t_{2}^{\ell(\Box)};p)_{\infty}}{(pq^{a(\Box)}t^{\ell(\Box)+1};p)_{\infty}},\qquad N^{\prime}_{\lambda}(p)=\prod_{\Box\in\lambda}\frac{(pt_{1}^{a(\Box)}t_{2}^{-\ell(\Box)-1};p)_{\infty}}{(pq^{-a(\Box)-1}t^{-\ell(\Box)};p)_{\infty}}.$

### v2

@@ Acknowledgements
$\Phi^{D}(u):{\mathcal{F}}^{(1,N)}_{v}\to{\mathcal{F}}^{(0,-1)}_{u}\widetilde{\otimes}{\mathcal{F}}^{(1,N+1)}_{-v/u}$,
$\Psi^{D*}(u)\ :\ {\mathcal{F}}^{(1,N)}_{v}\widetilde{\otimes}{\mathcal{F}}^{(0,-1)}_{u}\to{\mathcal{F}}^{(1,N-1)}_{-vu}$
@@
$x\Psi^{D*}(v)=\Psi^{D*}(v)\Delta^{D}(x)\qquad\qquad\forall x\in{\mathcal{U}}.$ (A.2)
@@
$\Phi^{D}(u)\xi=\sum_{\lambda\in{\mathcal{P}}^{+}}{|\lambda\rangle}^{\prime}_{u}\widetilde{\otimes}\Phi^{D}_{\lambda}(u)\xi,$ (A.3)
$\Psi^{D*}_{\lambda}(u)\xi=\Psi^{D*}(u)\left(\xi\widetilde{\otimes}{|\lambda\rangle}_{u}\right)\quad\forall\xi\in{\mathcal{F}}^{(1,N)}_{v}$, (A.4)
@@
$\hskip 5.69054pt{\Phi}^{D}_{\lambda}(u)=\frac{t_{1}^{-n(\lambda^{\prime})-|\lambda|}N_{\lambda}(p)}{c_{\lambda}}\hskip-2.84526pt:{\widetilde{\Phi}}_{\emptyset}(u)\prod_{\square\in\lambda}{x}^{-}(\varphi^{\lambda}_{\square}u):\hskip-2.84526pt(-u)^{-\alpha}e^{-\Lambda_{0}},$ (A.8)
$\hskip 5.69054pt{\Psi}^{D*}_{\lambda}(u)=\frac{t_{1}^{-n(\lambda^{\prime})}}{c_{\lambda}N_{\lambda}^{\prime}(p^{*})}\ :\widetilde{\Psi}^{*}_{\emptyset}(u)\prod_{\square\in\lambda}{x}^{+}(\varphi^{\lambda}_{\square}u):(-u)^{\alpha}e^{\Lambda_{0}},$ (A.9)
@@
$\hskip-25.60747ptN_{\lambda}(p)=\prod_{\Box\in\lambda}\frac{(pt_{1}^{-a(\Box)-1}t_{2}^{\ell(\Box)};p)_{\infty}}{(pq^{a(\Box)}t^{\ell(\Box)+1};p)_{\infty}},\ N^{\prime}_{\lambda}(p)=\prod_{\Box\in\lambda}\frac{(pt_{1}^{a(\Box)}t_{2}^{-\ell(\Box)-1};p)_{\infty}}{(pq^{-a(\Box)-1}t^{-\ell(\Box)};p)_{\infty}}.$